# HiMaCon: Discovering Hierarchical Manipulation Concepts from Unlabeled Multi-Modal Data

### v1

@@ Abstract
Effective generalization in robotic manipulation requires representations that capture invariant patterns of interaction across environments and tasks. We present a self-supervised framework for learning hierarchical manipulation concepts that encode these invariant patterns through cross-modal sensory correlations and multi-level temporal abstractions without requiring human annotation. Our approach combines a cross-modal correlation network that identifies persistent patterns across sensory modalities with a multi-horizon predictor that organizes representations hierarchically across temporal scales. Manipulation concepts learned through this dual structure enable policies to focus on transferable relational patterns while maintaining awareness of both immediate actions and longer-term goals. Empirical evaluation across simulated benchmarks and real-world deployments demonstrates significant performance improvements with our concept-enhanced policies. Analysis reveals that the learned concepts resemble human-interpretable manipulation primitives despite receiving no semantic supervision. This work advances both the understanding of representation learning for manipulation and provides a practical approach to enhancing robotic performance in complex scenarios.
@@ 1 Introduction
Figure 1: Manipulation concepts enhance generalization. Top: Training data with cups and containers without barriers. Middle: Without manipulation concepts, policies fail when encountering barriers. Bottom: With our concept enhancement, policies adapt by recognizing functional relationships and adjusting execution accordingly.
Enabling robots to perform manipulation tasks in diverse, unstructured environments remains a fundamental challenge in robotics. Despite advances in policy learning and sophisticated architectures (black2024pi_0, ; oxe, ; intelligence2025pi, ; kim2024openvla, ), current approaches often fail when encountering unexpected variations or novel scenarios. As illustrated in Fig. 1, a policy trained to place cups into containers may succeed in familiar settings but fail when encountering unexpected barriers – revealing a critical generalization gap that would limit real-world deployment.
We propose that addressing this generalization challenge requires learning transferable manipulation concepts – hierarchically structured abstractions that capture the fundamental patterns underlying manipulation. These concepts bridge the gap between low-level actions and high-level goals, enabling more robust generalization across environmental variations. For example, the concept of “placing an object inside a container” encompasses invariant relational patterns that persist whether the container has barriers or not, allowing a robot to adapt its specific actions while maintaining the core manipulation strategy.
@@ 2 Related Work
Representation Learning in Robotics. Self-supervised representation learning has emerged as a powerful approach for extracting meaningful patterns from robotic data without manual annotation (eisner2022flowbot3d, ; li2023manipllm, ; mo2019partnet, ). Initial efforts focused on single-modality approaches, with separate research streams developing for vision-based (bruce2024genie, ; dasari2023unbiased, ; seo2023masked, ; xu2023xskill, ; ye2024latent, ) and proprioception-based (lee2024behavior, ; mete2024quest, ; pertsch2025fast, ; shafiullah2022behavior, ) representation learning. Recent work has increasingly recognized the importance of multi-modal integration, combining vision with language (karamcheti2023language, ; majumdar2023we, ; nair2022r3m, ; schmidtlearning, ; wu2023unleashing, ), proprioception with vision (radosavovic2023robot, ; wangscaling, ; yang2023polybot, ), or even richer modality combinations (bonatti2023pact, ; shah2023mutex, ; zeng2024learning, ).
@@
Concept-Guided Robotic Policies. Concept-guided approaches enhance robotic policy performance by leveraging intermediate representations that bridge perception and action. These methods generally fall into two categories. Two-module frameworks (belkhale2024rt, ; bu2024towards, ; wangscaling, ; ye2024latent, ) employ a dual-model architecture where one component extracts high-level task concepts while another generates the corresponding actions. While effective, these approaches often require specialized architectural designs that limit their applicability across different policy classes.
In contrast, joint prediction approaches (guo2024prediction, ; jia2024chainofthought, ; shi2023waypoint, ; tian2025predictive, ; yang2022chain, ) integrate concept guidance by training policies to simultaneously predict both concepts and actions. This creates an implicit information pathway where concept understanding regularizes action generation. Our work adopts this more flexible approach, enabling seamless integration with diverse policy architectures while maintaining the interpretability benefits of explicit concept representations.
@@ 3 Method
Figure 2: The proposed self-supervised manipulation concept discovery and policy enhancement. Stage 1: The concept encoder ($\mathcal{E}$) processes multi-modal robot demonstrations to extract manipulation concept latents. These latents are refined through two complementary objectives: (1) the Cross-Modal Correlation Network ($\mathcal{C}$) employs a mask-and-predict strategy to capture persistent patterns across sensing modalities (Sec. 3.2); (2) the Multi-Horizon Future Predictor ($\mathcal{F}$) enables concept latents to organize hierarchically into short-horizon and long-horizon sub-goals based on coherence thresholds ($\epsilon$) (Sec. 3.3). Stage 2: The learned concepts are integrated into policy learning through a backbone network ($\pi_{h}$) with concept ($\pi_{z}$) and action ($\pi_{a}$) prediction heads, regularizing action generation with structured manipulation knowledge (Eq. 9).
We aim to encode robotic manipulation demonstrations into latent representations that capture task-induced patterns in multi-modal sensory-motor data. These representations should naturally cluster according to functional sub-goals, providing insights into manipulation objectives and enhancing policy learning. We term these clusters manipulation concepts – each representing action sequences targeting specific sub-goals – and call the learning process manipulation concept discovery.
@@ 3.1 Problem Setup & Manipulation Concept Encoder
Given a dataset $D=\{\tau_{i}\}_{i=1}^{N}$ with $N$ manipulation trajectories, each $\tau_{i}=\{(\mathbf{o}^{t}_{i},a^{t}_{i})\}_{t=1}^{T_{i}}$ contains observations $\mathbf{o}^{t}_{i}$ and actions $a^{t}_{i}$ at time $t$. For $M$ modalities, $\mathbf{o}^{t}_{i}=\{o^{1,t}_{i},o^{2,t}_{i},\dots,o^{M,t}_{i}\}$, where $o^{m,t}_{i}$ is the observation of the modality $m$. We denote $\mathbf{o}^{S,t}_{i}=\{o^{m,t}_{i}\mid m\in S\}$ as observations of modalities in $S\subseteq[M]=\{1,2,\cdots,M\}$. We treat observations of both the same sensory modes and different modes as distinct modalities, as they are functionally different modalities in terms of complementary information.
@@ 3.2 Capturing Multi-Modal Correlations
To learn manipulation concepts that capture cross-modal correlations, we propose maximizing mutual information – a metric capable of modeling diverse correlations – between observations from different modalities, conditioned on the associated manipulation concept. Specifically, we maximize the conditional mutual information over bipartitions of modality observations:
@@ 3.3 Representing Multi-Horizon Sub-Goals
where $K(\mathbf{z}_{i};\epsilon)$ is the number of clusters determined by $\epsilon$, and increasing $\epsilon$ yields sub-processes spanning from short-horizon to long-horizon. Please see Algorithm 1 for more details.
@@ 3.4 Enhancing Imitation Learning with Manipulation Concepts
Explicitly, with manipulation concepts $\mathbf{z}_{i}$ generated by encoder $\mathcal{E}$, we augment imitation learning by training policies to predict both ground-truth actions and corresponding concepts (jia2024chain, ; yang2022chain, ; zhao2025cot, ). This approach uses concept prediction as a regularization that guides the policy to encode conceptual understanding alongside action planning:
@@ 4.1 Experimental Setup
  • InfoCon (liuinfocon, ): VQ-VAE type of method for single-hierarchy concept discovery.
@@ 4.2 Evaluating Policy Performance with Learned Manipulation Concepts
Performance on Original Training Tasks
We first evaluate our concept discovery method on the same tasks used for concept training. As shown in the L90-90 results (Tab. 1), our approach consistently outperforms all baselines with both policy architectures. The performance gap between our method and Next/InfoCon demonstrates the importance of multi-hierarchical sub-goal modeling, while improvements over All highlight the value of explicitly capturing cross-modal correlations. Our method also surpasses DecisionNCE variants despite not requiring language supervision, validating the effectiveness of our self-supervised objectives.
Transfer to Long-Horizon Tasks
To evaluate concept transferability to more complex compositions, we apply concept encoders trained on LIBERO-90 directly to LIBERO-LONG demonstrations featuring novel long-horizon tasks. The L90-L results show our method maintains its performance advantage in this challenging transfer setting. This demonstrates that our approach learns manipulation concepts that effectively decompose hierarchical tasks, enabling policies to better handle novel complex task compositions requiring sequential execution of multiple sub-goals.
Generalization to Novel Environments
We further test generalization by applying concept encoders trained on LIBERO-90 directly to LIBERO-GOAL demonstrations featuring unseen environments and tasks. The L90-G results show our method continues to outperform all baselines in this challenging scenario. This indicates our approach discovers fundamental manipulation primitives that transfer effectively across environmental variations.
Impact of Multi-Modal Observations
Our ablation study (Tab. 2) shows that performance consistently improves as more modalities are incorporated. The most significant drops occur when removing proprioceptive information, highlighting its importance for grounding visual observations with physical interaction states and confirming the value of our cross-modal correlation approach.
@@ 4.3 Analyzing Manipulation Concept Properties
To verify our Cross-Modal Correlation Network’s effectiveness, we measure mutual information between modalities conditioned on concept latents (based on (belghazi2018mutual, ), detailed in Sec. A.4). Tab. 4.3 shows our approach achieves higher conditional mutual information than the All baseline. This confirms our mask-and-predict strategy enables the concept encoder to capture persistent cross-modal patterns that generalize across different objects and contexts, providing a robust representational basis for manipulation policies.
Table 3: Conditional mutual information between modality pairs. Values conditioned on manipulation concept latents from our method versus the All baseline that does not model cross-modal correlations. A: agentview, H: eye-in-hand vision, P: proprioceptives.
@@
We evaluate whether our concepts align with human-understandable semantics by grouping latents from different demonstrations based on human-identified sub-goals (e.g., “reach for drawer,” “grasp handle,” “pull drawer”) and computing similarities between these groupings using Eq. 10:
@@
Varying the coherence threshold $\epsilon$ in Eq. 4 reveals the hierarchical organization of our learned concepts. Fig. 6 (Sec. C.5) shows larger $\epsilon$ values identify coarse-grained phases, while smaller values capture fine-grained actions. This emergent hierarchy enables policies to simultaneously reason about immediate actions and longer-term goals without explicit hierarchical supervision, contributing to improved performance on complex sequential tasks that require coordinated execution across multiple temporal scales.
Figure 3: Semantic alignment of learned concepts. Cosine similarity between concept latents grouped by human-defined sub-goals. Strong diagonal patterns demonstrate that our approach discovers concepts that exhibit clustering patterns corresponding to meaningful manipulation primitives.
@@ 4.4 Real-World Validation
To study generalization capabilities, we deploy concept-enhanced policies on a Mobile ALOHA robot (fu2024mobile, ) in “cleaning cup” tasks (Fig. 4). Training data includes only simple container arrangements with consistent color pairings. We test on six increasingly challenging variations: (1) Novel Placements: Cups and containers in unseen arrangements; (2) Color Composition: Altered cup-container color pairings; (3) Novel Objects: Entirely unseen containers, cups, and plates; (4) Obstacles: Objects between the robot and the cups obstructing vision; (5) Barriers: Internal dividers within containers impeding placement; (6) Grasping Together: Two adjacent cups requiring simultaneous grasp.
@@
  • Relational focus: Concept-enhanced policies emphasize transferable relational patterns (e.g., “object inside container”), enabled by our approach’s emphasis on cross-modal correlations.
  • Hierarchical awareness: Concept-enhanced policies demonstrate more effective recovery behaviors after failed attempts, suggesting awareness of incomplete sub-goals, while baseline policies more often fail to complete task stages (e.g., hovering near containers without placing cups).
@@ A.1 Manipulation concept discovery (Ours)
  • A granularity parameter $\epsilon\in[0,1]$ is incorporated by processing a uniformly sampled $\epsilon$ from $[0,1]$ using a 2-layer MLP with a hidden feature dimension of 1024. This is transformed into a 256-dimensional embedding, which is added to every time step of the input feature sequence to the transformer in $\mathcal{F}$.

### v2

@@ Abstract
Effective generalization in robotic manipulation requires representations that capture invariant patterns of interaction across environments and tasks. We present a self-supervised framework for learning hierarchical manipulation concepts that encode these invariant patterns through cross-modal sensory correlations and multi-level temporal abstractions without requiring human annotation. Our approach combines a cross-modal correlation network that identifies persistent patterns across sensory modalities with a multi-horizon predictor that organizes representations hierarchically across temporal scales. Manipulation concepts learned through this dual structure enable policies to focus on transferable relational patterns while maintaining awareness of both immediate actions and longer-term goals. Empirical evaluation across simulated benchmarks and real-world deployments demonstrates significant performance improvements with our concept-enhanced policies. Analysis reveals that the learned concepts resemble human-interpretable manipulation primitives despite receiving no semantic supervision. This work advances both the understanding of representation learning for manipulation and provides a practical approach to enhancing robotic performance in complex scenarios. Code is available at: https://github.com/zrllrz/HiMaCon.
@@ 1 Introduction
Figure 1: Manipulation concepts enhance generalization. Top: Training data with cups and containers without barriers. Middle: Without manipulation concepts, policies fail when encountering barriers. Bottom: With our concept enhancement, policies adapt accordingly.
Robot manipulation in diverse, unstructured environments remains a fundamental challenge. Despite advances in policy learning and architectures (black2024pi_0, ; oxe, ; intelligence2025pi, ; kim2024openvla, ), current approaches often fail when encountering unexpected variations or novel scenarios. As illustrated in Fig. 1, a policy trained to place cups into containers may succeed in familiar settings but fail when encountering unexpected barriers—revealing a critical generalization gap limiting real-world deployment.
We propose that addressing this challenge requires learning transferable manipulation concepts—hierarchical abstractions capturing fundamental manipulation patterns. These concepts connect low-level actions to high-level goals, enabling robust generalization. For example, the concept of “placing an object inside a container” encompasses invariant relational patterns that persist whether the container has barriers or not, allowing adaptation while maintaining core manipulation strategy.
@@ 2 Related Work
Representation Learning in Robotics
Self-supervised representation learning has emerged as a powerful approach for extracting meaningful skills (liang2024skilldiffuser, ; liu2025one, ; rholanguage, ) from robotic data, avoiding the need for manual annotation in methods such as (eisner2022flowbot3d, ; li2023manipllm, ; mao2023learning, ; mo2019partnet, ). Initial efforts explored single-modality approaches for vision-based (bruce2024genie, ; dasari2023unbiased, ; seo2023masked, ; xu2023xskill, ; ye2024latent, ) and proprioception-based (lee2024behavior, ; mete2024quest, ; pertsch2025fast, ; shafiullah2022behavior, ) representation learning. Recent work integrates multiple modalities, combining vision with language (karamcheti2023language, ; majumdar2023we, ; nair2022r3m, ; schmidtlearning, ; wu2023unleashing, ), proprioception with vision (radosavovic2023robot, ; wangscaling, ; yang2023polybot, ), even richer (bonatti2023pact, ; shah2023mutex, ; zeng2024learning, ).
@@
Concept-Guided Robotic Policies
Concept-guided approaches enhance robotic policy performance by leveraging intermediate representations that bridge perception and action. These methods generally fall into two categories. First, two-module frameworks (belkhale2024rt, ; bu2024towards, ; liang2024skilldiffuser, ; wangscaling, ; ye2024latent, ; zhouautocgp, ) employ a dual-model architecture where one component extracts high-level task concepts while another generates the corresponding actions. While effective, these approaches often require specialized architectural designs that limit their applicability across different policy classes.
Second, in contrast, joint prediction approaches (guo2024prediction, ; jia2024chainofthought, ; shi2023waypoint, ; tian2025predictive, ; yang2022chain, ) integrate concept guidance by training policies to simultaneously predict both concepts and actions. This creates an implicit information pathway where concept understanding regularizes action generation. Our work adopts this more flexible approach, enabling seamless integration with diverse policy architectures while maintaining the interpretability benefits of explicit concept representations.
@@ 3 Method
Figure 2: The proposed self-supervised manipulation concept discovery and policy enhancement. Stage 1: The concept encoder ($\mathcal{E}$) processes multi-modal robot demonstrations to extract concept latents. These latents are refined through two objectives: (1) the Cross-Modal Correlation Network ($\mathcal{C}$) employs a mask-and-predict strategy to capture persistent patterns across sensing modalities (Sec. 3.2); (2) the Multi-Horizon Future Predictor ($\mathcal{F}$) enables concept latents to organize hierarchically into multi-horizon sub-goals based on coherence thresholds ($\epsilon$) (Sec. 3.3). Stage 2: The learned concepts are integrated into policy learning through a backbone network ($\pi_{h}$) with concept ($\pi_{z}$) and action ($\pi_{a}$) prediction heads, regularizing action generation with structured manipulation knowledge (Eq. 9).
We aim to encode robotic manipulation demonstrations into latent representations that capture task-induced patterns in multi-modal sensory-motor data. These representations should naturally cluster according to functional sub-goals, providing insights into manipulation objectives and enhancing policy learning. We term these clusters manipulation concepts—each representing action sequences targeting specific sub-goals—and call the learning process manipulation concept discovery.
@@ 3.1 Problem Setup & Manipulation Concept Encoder
Given a dataset $D=\{\tau_{i}\}_{i=1}^{N}$ of $N$ manipulation trajectories, each $\tau_{i}=\{(\mathbf{o}^{t}_{i},a^{t}_{i})\}_{t=1}^{T_{i}}$ contains observations $\mathbf{o}^{t}_{i}$ and actions $a^{t}_{i}$ at time $t$. For $M$ modalities, $\mathbf{o}^{t}_{i}=\{o^{1,t}_{i},o^{2,t}_{i},\dots,o^{M,t}_{i}\}$, where $o^{m,t}_{i}$ is the observation of the modality $m$. We denote $\mathbf{o}^{S,t}_{i}=\{o^{m,t}_{i}\mid m\in S\}$ as observations of modalities in $S\subseteq[M]=\{1,2,\cdots,M\}$. We treat observations of both the same sensory modes (e.g., multiple views) and different modes as distinct modalities, as they are functionally different modalities in terms of complementary information.
@@ 3.2 Capturing Multi-Modal Correlations
To learn manipulation concepts that capture cross-modal correlations, we propose maximizing mutual information—a metric capable of modeling diverse correlations—between observations from different modalities, conditioned on the associated manipulation concept. Specifically, we maximize the conditional mutual information over bipartitions of modality observations:
@@ 3.3 Representing Multi-Horizon Sub-Goals
where $K(\mathbf{z}_{i};\epsilon)$ is the number of clusters determined by $\epsilon$, and increasing $\epsilon$ yields sub-processes spanning from short-horizon to long-horizon. Please see Alg. 1 for more details.
@@ 3.4 Enhancing Imitation Learning with Manipulation Concepts
Specifically, with manipulation concepts $\mathbf{z}_{i}$ generated by encoder $\mathcal{E}$, we augment imitation learning by training policies to predict both ground-truth actions and corresponding concepts (jia2024chain, ; yang2022chain, ; zhao2025cot, ). This approach uses concept prediction as a regularization that guides the policy to encode conceptual understanding alongside action planning:
@@ 4.1 Experimental Setup
  • InfoCon (liuinfocon, ): A VQ-VAE type of method for single-hierarchy concept discovery.
@@ 4.2 Evaluating Policy Performance with Learned Manipulation Concepts
  • Performance on Original Training Tasks  We first evaluate our concept discovery method on the same tasks used for concept training. As shown in the L90-90 results (Tab. 1), our approach consistently outperforms all baselines with both policy architectures. The performance gap between our method and Next/InfoCon demonstrates the importance of multi-hierarchical sub-goal modeling, while improvements over All highlight the value of explicitly capturing cross-modal correlations. Our method also surpasses DecisionNCE variants despite not requiring language supervision, validating the effectiveness of our self-supervised objectives.
  • Transfer to Long-Horizon Tasks  To evaluate concept transferability to more complex compositions, we apply concept encoders trained on LIBERO-90 directly to LIBERO-LONG demonstrations featuring novel long-horizon tasks. The L90-L results show our method maintains its performance advantage in this challenging transfer setting. This demonstrates that our approach learns manipulation concepts that effectively decompose hierarchical tasks, enabling policies to better handle novel complex task compositions requiring sequential execution of multiple sub-goals.
  • Generalization to Novel Environments  We further test generalization by applying concept encoders trained on LIBERO-90 directly to LIBERO-GOAL demonstrations featuring unseen environments and tasks. The L90-G results show our method continues to outperform all baselines in this challenging scenario. This indicates our approach discovers fundamental manipulation primitives that transfer effectively across environmental variations.
  • Impact of Multi-Modal Observations  Our ablation study (Tab. 2) shows that performance consistently improves as more modalities are incorporated. The most significant drops occur when removing proprioceptive information, highlighting its importance for grounding visual observations with physical interaction states and confirming the value of our cross-modal correlation approach.
@@ 4.3 Analyzing Manipulation Concept Properties
Table 3: Conditional mutual information between modality pairs. Values conditioned on concept latents from our method versus the All baseline that does not model cross-modal correlations. A: agentview, H: eye-in-hand vision, P: proprioception.
@@
To verify our Cross-Modal Correlation Network’s effectiveness (Sec. 3.2), we measure mutual information between modalities conditioned on concept latents (Sec. A.4). Tab. 4.3 shows that our approach achieves higher conditional mutual information than the All baseline. This confirms that our mask-and-predict strategy enables the concept encoder to capture persistent cross-modal patterns that generalize across different objects and contexts, providing a robust representational basis for policies.
@@
We evaluate whether our concepts align with human-understandable semantics by grouping latents from different demonstrations based on human-identified sub-goals and computing similarities between these groupings:
@@
Varying the coherence threshold $\epsilon$ in Eq. 4 reveals the hierarchical organization of our learned concepts. Fig. 3 (and Sec. C.5) shows larger $\epsilon$ values identify coarse-grained phases, while smaller values capture fine-grained actions. This emergent hierarchy enables policies to simultaneously reason about immediate actions and longer-term goals without explicit hierarchical supervision, contributing to improved performance on complex sequential tasks that require coordinated execution across multiple temporal scales.
@@ 4.4 Real-World Validation
To study generalization capabilities, we deploy concept-enhanced policies on a Mobile ALOHA robot (fu2024mobile, ) in “cleaning cup” tasks (Fig. 5). Training data includes only simple container arrangements with consistent color pairings.
Figure 4: Semantic alignment of learned concepts. Cosine similarity between concept latents grouped by human-defined sub-goals. Diagonal patterns demonstrate that our approach discovers concepts that exhibit clustering patterns corresponding to meaningful manipulation primitives.
We test on six increasingly challenging variations: (1) Novel Placements: Cups and containers in unseen arrangements; (2) Color Composition: Altered cup-container color pairings; (3) Novel Objects: Entirely unseen containers, cups, and plates; (4) Obstacles: Objects between the robot and the cups obstructing vision; (5) Barriers: Internal dividers within containers impeding placement; (6) Grasping Together: Two adjacent cups requiring simultaneous grasp.
@@
1. Relational focus: Concept-enhanced policies prioritize transferable relational patterns (e.g., “object inside container”) over surface features. Our cross-modal correlation learning (Sec. 3.2) enables this capability by identifying patterns that remain invariant across modalities. This relational emphasis explains the stronger performance on scenarios that alter visual appearance while preserving task structure. For instance, while Novel Placement tests spatial variation alone, the other protocols introduce substantial visual perturbations (different colors, objects, or occlusions) that shift the appearance distribution. The consistent performance gains across these visually diverse scenarios (Tab. 4.4) suggest that the learned concepts successfully capture the underlying relational invariant —placing cups into containers —rather than memorizing superficial visual patterns.
2. Hierarchical awareness: Concept-enhanced policies exhibit more systematic failure recovery than baselines, suggesting better tracking of sub-goal completion. Baseline failures frequently exhibit premature task abandonment: the robot moves toward containers without having grasped objects, or hovers near placement locations without executing placement. In contrast, when concept-enhanced policies fail initial grasp attempts, they consistently retry grasping (typically 2-3 attempts) before proceeding, demonstrating recognition of incomplete sub-goals. Although these recoveries ultimately fail due to time limits or object displacement, they reveal structured task progression rather than blind action execution.
@@ 5 Discussion
Acknowledgments and Disclosure of Funding
This work is supported by the Early Career Scheme of the Research Grants Council (RGC) grant # 27207224, the HKU-100 Award, a donation from the Musketeers Foundation, in part by the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust, and DAMO Academy through the Alibaba Innovative Research Program.
@@ A.1 Manipulation concept discovery (Ours)
  • To incorporate the granularity parameter $\epsilon\in[0,1]$, we discretize the continuous range into 1000 uniform bins $\{0.000,0.001,\ldots,0.999\}$ and learn a corresponding VQ-VAE codebook [59] with 1000 entries, each represented as a 256-dimensional embedding vector. In each transformer block, the feed-forward layer receives the concatenation of the attention output and the embedding corresponding to the sampled $\epsilon$ value.